OpenStudy (anonymous):

Is this a sum of squares? Can someone factor this: g(x) = x^2+92

OpenStudy (mathstudent55):

Is it \(92\) or \(9^2\)? Depending on what it is, the answer is either no or yes.

OpenStudy (mathstudent55):

\(a^2 - b^2 = (a + b)(a - b)\) \(a^2 + b^2 = a^2 - (-b^2) = (a + bi)(a - bi)\)

OpenStudy (anonymous):

it's 9^2 sorry @mathstudent55

OpenStudy (anonymous):

but actually could you explain to me Function: g(x) = x^2 + 2^2 can't be factored? I think it's something with prime numbers but I can't remember

OpenStudy (anonymous):

@mathstudent55 you there?

OpenStudy (anonymous):

@FortyTheRapper could you help explain for me

OpenStudy (fortytherapper):

Is that x^2+4 basically?

OpenStudy (anonymous):

i think so, because 2 squared is 4

OpenStudy (fortytherapper):

It's pretty much because of that + sign. In order to factor these types, they would need to be a minus sign

OpenStudy (anonymous):

How would I put that in writing as to why it can't be factored. Would I say because the equation is prime?

OpenStudy (mathstudent55):

A sum of squares cannot be factored. A difference of squares can be factored.

OpenStudy (fortytherapper):

^ I like that one

OpenStudy (whpalmer4):

A sum of squares can be factored, if you are willing to use complex numbers. \[x^2+9^2 = x^2-(-1)(9^2) = x^2-(i^2)(9^2) = (x+3i)(x-3i)\]where \(i=\sqrt{-1}\)

OpenStudy (anonymous):

oh okay!

OpenStudy (whpalmer4):

sorry, that should be \[(x+9i)(x-9i)\]at the end...

OpenStudy (anonymous):

what would you say the method is to factor a difference of squares?

OpenStudy (whpalmer4):

recognition of a pattern. Anytime I see <something>^2 - <something else>, the little bell goes off and I try to figure out how I can express <something else> as a square.

OpenStudy (anonymous):

what do you mean?

OpenStudy (whpalmer4):

\[x^2-9\]well, obviously \(x^2\) is something squared...so now I look at the \(9\) and say "how do I write \(9\) as some other quantity squared?" well, \[9=a^2\]\[\sqrt{9} = a\]\[3=a\]so \[x^2-9 = x^2-a^2 = x^2-(3)^2\]so I can immediately write that factored as \[x^2-9 = (x+3)(x-3)\]

OpenStudy (whpalmer4):

try factoring this: \[16x^2-25\]

OpenStudy (anonymous):

Do I start looking for the GCF?

OpenStudy (whpalmer4):

there is no GCF in this case...just look at it and tell me if it is a difference of squares.

OpenStudy (anonymous):

No, I think

OpenStudy (anonymous):

Oh wait, I think it is

OpenStudy (anonymous):

(4x - 5) (4x +5)

OpenStudy (anonymous):

Or (4x + 5) (4x - 5) does the matter the order in which the + or - is placed?

OpenStudy (whpalmer4):

nope, doesn't matter. multiplication is commutative. \[(4x+5)(4x-5) = (4x-5)(4x+5)\] Here's a slightly trickier one: \[9x^2y^6-100z^4\]

OpenStudy (anonymous):

Oh wow, um, (3xy^3 - 10z^2) (3xy^3 - 10z^2) ?

OpenStudy (anonymous):

Could you help me understand a perfect square trinomial?

OpenStudy (whpalmer4):

yes, you got it it right, well, almost right...doesn't one of those terms need a + instead of a -?

OpenStudy (anonymous):

Oh Right, I forgot that part! It needs it to make the other question have the -100

OpenStudy (whpalmer4):

a perfect square trinomial is one obtained by squaring a binomial. \[(x+3)^2 = (x+3)(x+3) = x*x + x*3 + 3*x + 3*3 = x^2 + 6x + 9\]That is a perfect square trinomial. We can recognize a PST by a couple of traits: 1) 1st term is a positive perfect square 2) 3rd term is a positive perfect square 3) 2nd term is twice the product of the square root of the first term and square root of the third term Let's check our prospective PST: 1) 1st term is positive perfect square. \(x^2\) is positive and a perfect square. 2) 3rd term is positive perfect square. \(9\) is positive and a perfect square. 3) 2nd term is twice product of square root of 1st term and square root of 3rd term.\[6x = 2\sqrt{x^2}\sqrt{9} = 2*x*3 = 6x\checkmark\]

OpenStudy (whpalmer4):

\[x^2-6x+9\]Is that a PST?

OpenStudy (anonymous):

No because the 1st and 3rd term does not result in the 2nd term right?

OpenStudy (whpalmer4):

uh, well, I must have misstated the rule :-) \[(x-3)^2=(x-3)(x-3) = x^2 - 3x -3x +9 = x^2 -6x + 9\] So it IS a PST. 2nd term (ignoring sign) must be equal to 2*sqrt(1st term)*sqrt(3rd term)

OpenStudy (anonymous):

Oh okay, so in other words: The function g(x) is a perfect square trinomial. Function: g(x)=(x-3)^2 Factored: g(x)=x^2-6x+9 That's how my school teaches it

OpenStudy (anonymous):

Also do you know what it means if a function could only have a GCF factored out of it? Would it almost be like the difference of squares?

OpenStudy (whpalmer4):

uh, factored would be \[g(x) = (x-3)(x-3)\] \[x^3+3x^2+7x\]can only have a GCF of \(x\) factored out of it: \[x(x^2+3x+7)\]but that doesn't imply anything about what's left...

OpenStudy (anonymous):

So it needs the gcf to be factored out in order to get a result?

OpenStudy (whpalmer4):

if you can only factor out a GCF from the function, that's all the factoring you can do...

OpenStudy (anonymous):

Oh! Okay, could you explain the rest of what would be implied about what's left?

OpenStudy (whpalmer4):

That simply means that what is left is not factorable. For example, you can't factor \[x^2+3x+7\]if you stick to rational numbers. I can multiply that polynomial by anything I want, and that will give me another polynomial where I can factor out as a GCF whatever I multiplied by, but what's left is not factorable. I could also multiply a polynomial that IS factorable by a GCF: \[x^4-9x^2\] Can you factor that?

OpenStudy (anonymous):

Oh, that makes sense I think. And, x^2(x^2 - 9) ?

OpenStudy (whpalmer4):

factor it all the way out...

OpenStudy (anonymous):

There's more to that? Um (x^2-3) (x^1+9) I'm not doing this right am I ahah?

OpenStudy (whpalmer4):

\[x^2(x^2-9) = x*x*(x-3)(x+3)\]

OpenStudy (whpalmer4):

\[x^2-9\]is a difference of squares, right?

OpenStudy (anonymous):

no, but what could it be?

OpenStudy (whpalmer4):

what do you mean, no? \[(x+3)(x-3) = x*x -3x+3x-9 = x^2-9\]Is that not a difference of squares?

OpenStudy (anonymous):

wouldn't it have to be x^4 - 9 in the end?

OpenStudy (anonymous):

I mean x^2 - 9 is a difference of squares but that doesn't go back to x^4 - 9. Am I even making sense right now? Am I right or wrong

OpenStudy (mathstudent55):

\(x^4 - 9\) can be rewritten as \((x^2)^2 - 3^2\) where you see clearly it is the difference of two squares. It is the square of \(x^2\) minus the square of \(3\). \(x^4 - 9 = (x^2 + 3)(x^2 - 3) \)

OpenStudy (whpalmer4):

\[x^4-9x^2\]can be factored as follows both terms have a GCF of \(x^2\): \[x^4-9x^2=x^2*x^2-9*x^2=x^2(x^2-9)\]the stuff in parentheses is a difference of squares\[x^2(x^-9)=x^(x^2-3^2)=x^2(x-3)(x+3)\] \[x^4-9\]is also a difference of squares:\[(x^2)^2-3^2=(x^2-3)(x^2+3)\]if we had gone with a different exampl,\[x^4-81 = (x^2)^2-9^2=(x^2-9)(x^2+9)\]but \(x^2-9\) is a difference of squares so that factors further into\[x^4-81=(x^2-9)(x^2+9)=(x-3)(x+3)(x^2+9)\]